OpenStudy (anonymous):

Give the standard form of the equation. Hyperbola with vertex at (0, 4), focus at (0, 5), and center at (0, 1)

OpenStudy (mertsj):

a^2+b^2=c^2 c=5 and a=4

OpenStudy (anonymous):

can u give it to me in this form.

OpenStudy (mertsj):

That's why I showed you how to find a, b, and c so you could write it in that form.

OpenStudy (mertsj):

\[\frac{(y-k)^2}{b^2}-\frac{(x-h)^2}{a^2}=1\]

OpenStudy (mertsj):

\[\frac{(y-1)^2)}{16}-\frac{(x-0)^2}{9}=1\]